OpenStudy (zab505):

Suppose that a customer service center claims that the telephone calls it receives last 112 seconds on average, with a standard deviation of 9 seconds. If you took a sample of 324 telephone calls, which of the following mean times would be within the 95% confidence interval? A. 113.5 seconds B. 114.5 seconds C. 112.5 seconds D. 115.5 seconds

OpenStudy (anonymous):

The confidence interval has the form \[\left(\bar{x}-Z_{\alpha/2}\frac{\sigma}{\sqrt n},~\bar{x}+Z_{\alpha/2}\frac{\sigma}{\sqrt n}\right)\] where \(\bar{x}\) is the sample mean (112 seconds), \(Z_{\alpha/2}\) is the cutoff for the \((1-\alpha)100\%\) confidence level, \(\sigma\) is the sample/population standard deviation (9 seconds), and \(n\) is the sample size (324 callers). Since the confidence level is 95%, you have \(Z_{\alpha/2}=1.96\). Construct the interval and check which of the listed values fall in its range.